The function Q (t)= 0.003t^2 - 0.625t + 25 represents the amount of energy in a battery after t minutes of use.
(a) State the amount of energy held by the battery immediately before it was used.
(b) Calculate the amount of energy available after 20 minutes.
(c) Given that Q (10) = 19.05, find the average amount of energy produced per minute for the interval 10 ≤ t ≤ 20
(d) Calculate the number of minutes it takes for the energy to reach zero.

Answer by htmentor(1343)

(a) At t=0, Q(0) = 25
(b) Q(20) = 0.003(20)^2 - 0.625(20) + 25 = 13.7
(c) Average amount of energy per minute over the interval [10,20] =
(13.7 - 19.05)/(20-10) = -0.535 per min. The average rate is negative, since
energy is expended.
(d) Setting Q = 0, we have 0.003t^2 - 0.625t + 25 = 0
Solve with the quadratic formula. The first zero crossing is at t = 53.99, ~54 min